# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

### v2

@@ 3.1 Preliminaries
Formally, for a document $\mathcal{D}$ composed of $N$ words, denoted as $\mathcal{D}=\{w_{1},w_{2},\ldots,w_{N}\}$, a set of pre-defined event types $\mathcal{E}$, a corresponding set of argument roles $\mathcal{R}_{e}$ associated with each event $e\in\mathcal{E}$, and the event trigger $t$ in $\mathcal{D}$. The goal is to detect every pair $(r,s)$ for every event $e$ within the document $\mathcal{D}$, where $r\in\mathcal{R}_{e}$ signifies a role associated with the event $e$, and $s$ denotes a contiguous text span within $\mathcal{D}$. In alignment with earlier research xu2022two ; ebner2020multi , we extract event arguments independently for each event within a document.
3.2 Enhanced Span-based Baseline
In this section, we improve existing span-based baselines ebner2020multi ; zhang2020two and introduce a new baseline for document-level EAE. We integrate other techniques to further improve the performance and the architecture is shown in Figure 3.
3.2.1 Encoder
Following xu2022two ; ebner2020multi , we divide a document into several instances, where each instance only contains one event. For each document instance $\mathcal{D}$, the target event $e$ and the given trigger word $t_{e}$, we wrap the event trigger with special tokens "*""""*"" * " and regard the embedding of the start "*""""*"" * " token as the trigger embedding, which is adapted from the entity marker technique shi2019simple ; soares2019matching . We also add the event type as additional information guidance and the final input sequence $S$ is obtained as follows:
$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}*\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]},$
where $[E_{e}]$ is the special event token [^1] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 . of event $e$. Next, we use a pre-trained transformer encoder to get the embedding of each token as follows:
$\mathbf{H}=\mathrm{Encoder}(\textsc{$S$}).$ (1)
@@ 3.2.1 Encoder
Next, we can get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start $[E_{e}]$ and the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, where $l_{w}$ represents the length of the word pieces list. Notably, when the input sequence is longer than the maximum token length, we employ a dynamic window approach zhou2021document ; zhang2021document to process the long sequence. We divide the sequence into overlapping windows and calculate the embeddings for each window. Next, we obtain the final representation by averaging the embeddings of tokens that overlap across various windows.
3.2.2 Candidate Spans Representation
Filtering Impossible Spans   To eliminate noise information from useless spans, we implement a filtering step to reduce the number of candidate spans. Specifically, we exclude argument-impossible spans, such as spans containing commas in the middle. For a document $\mathcal{D}$, candidate spans after argument-impossible exclusion can be derived as $\mathcal{S}=\{s_{i:j}\}$. This enhancement allows us to eliminate approximately 25% of candidate spans on average, enabling our model to focus on candidate spans that contain valuable information.
Candidate Spans Representation  For a candidate argument span ranging from $w_{i}$ to $w_{j}$, we first obtain its average pooling representation $\textbf{s}_{i:j}$ as follows:
$\textbf{s}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w},$ (2)
where $\mathbf{h}_{k}^{w}$ is the $k^{\mathrm{th}}$ token embedding from $\mathbf{H}^{w}$.
As we extract arguments at the span level, where the boundaries can be ambiguous, we employ fully connected neural networks to construct start and end representations as follows:
$\mathbf{H}^{start}=\mathbf{W}^{start}\mathbf{H},$ (3)
$\mathbf{H}^{end}=\mathbf{W}^{end}\mathbf{H},$ (4)
where $\mathbf{H}$ represents the hidden representation of the input sequence $S$, and $\mathbf{W}^{start},\mathbf{W}^{end}\in\mathbb{R}^{d\times d}$ are trainable parameters. Subsequently, the candidate span representation $\widetilde{\mathbf{s}}_{i:j}$ can be obtained as follows:
$\widetilde{\mathbf{s}}_{i:j}=\mathbf{W}^{s}[\mathbf{h}^{start}_{i};\textbf{s}_{i:j};\mathbf{h}^{end}_{j}],$ (5)
where $\mathbf{W}^{s}\in\mathbb{R}^{3d\times d}$ is the trainable parameter and $\mathbf{h}^{start}_{i}$ and $\mathbf{h}^{end}_{j}$ are the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ vector of $\mathbf{H}^{start}$ and $\mathbf{H}^{end}$. we establish the boundary loss for identifying the start and end positions in accordance with xu2022two :
$\begin{array}[]{c}\mathcal{L}_{b}=-\sum_{i=1}^{|\mathcal{D}|}[y_{i}^{s}\log P_{i}^{s}+(1-y_{i}^{s})\log(1-P_{i}^{s})\\ +y_{i}^{e}\log P_{i}^{e}+(1-y_{i}^{e})\log(1-P_{i}^{e})],\end{array}$ (6)
where $y_{i}^{s}$ and $y_{i}^{e}$ denote golden labels and $P_{i}^{s}=\mathrm{sigmoid}(\mathbf{W}_{1}\mathbf{h}^{start}_{i})$ and $P_{i}^{e}=\mathrm{sigmoid}(\mathbf{W}_{2}\mathbf{h}^{end}_{i})$ represent the probabilities of the word $w_{i}$ being predicted as the start or end word of a golden span.
3.2.3 Classification Module
Within each event $e$, for each candidate span $s_{i:j}$, we apply a comparison technique mou-etal-2016-natural ; zeng2020double to enhance its representation. In detail, we concatenate the span representation $\widetilde{\mathbf{s}}_{i:j}$ with the trigger representation $\mathbf{h}_{t}$, their absolute disparity $|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|$, element-wise product $\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j}$, the embedding of the event type $\mathbf{H}^{e}$ and the embedding of the span length $\mathbf{E}_{len}$. Finally, through a feed-forward network, we calculate the predicted probability $P(r_{s_{i:j}})$ of the candidate span $s_{i:j}$:
$\mathbf{I}_{i:j}=[\widetilde{\mathbf{s}}_{i:j};\mathbf{h}_{t};|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|;\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j};\mathbf{H}^{e};\mathbf{E}_{len}],$ (7)
$P(r_{s_{i:j}})=\mathrm{FFN}(\mathbf{I}_{i:j}).$ (8)
Considering the issue of imbalanced role distribution and the prevalence of negative samples among candidate arguments, we utilize focal loss 8237586 during the training process. Focal loss uses hyperparameters $\alpha$ and $\gamma$ to control the weighting and scaling of the loss function, which allows us to emphasize the importance of positive samples and makes the model focus more on useful information.
$\begin{array}[]{c}\mathcal{L}_{c}=-\sum_{s_{i:j}\in\mathcal{S}}\alpha{{[1-P(r_{s_{i:j}}=y_{s_{i:j}}}))]}^{\gamma}\\ \cdot\log P(r_{s_{i:j}}=y_{s_{i:j}}).\end{array}$ (9)

### v4

@@ 3.2 Task Formulation
Formally, for a document $\mathcal{D}$ composed of $N$ words, denoted as $\mathcal{D}=\{w_{1},w_{2},\ldots,w_{N}\}$, a set of pre-defined event types $\mathcal{E}$, a corresponding set of argument roles $\mathcal{R}_{e}$ associated with each event $e\in\mathcal{E}$, and the event trigger $t$ in $\mathcal{D}$. The goal is to detect every pair $(r,s)$ for every event $e$ within the document $\mathcal{D}$, where $r\in\mathcal{R}_{e}$ signifies a role associated with the event $e$, and $s$ denotes a contiguous text span within $\mathcal{D}$. In alignment with earlier research (Xu et al., 2022b; Ebner et al., 2020), we extract event arguments independently for each event within a document.
4 CARLG
In this section, we introduce CARLG framework, comprising Contextual Clues Aggregation (CCA) module which adaptively aggregates contextual clues and Role-based Latent Information Guidance (RLIG) module which effectively captures semantic correlations among roles. The core purpose of these two modules is to enhance the representation of the arguments to be predicted. Additionally, they can be applied for any approach that relies on argument representation, including span-based methods that make predictions using candidate span representation, and prompt-based methods that predict based on argument slot representation.
4.1 Contextual Clues Aggregation
Given that different events focus on different contextual information, the CCA module adaptively aggregates contextual information based on specific event feature to obtain a broader context representation. Since the event trigger is crucial for the arguments within an event, we adaptively aggregate context information based on each trigger-argument pair, denoted as $(t,a_{i})$ and $a_{i}$ is the $i^{th}$ argument. We leverage the attentive heads from the pre-trained transformer encoder, which enables the adoption of attentions from the pre-trained language model instead of training new attention layers. Figure 3 illustrates the process of the proposed CCA module.
Specifically, we encode the input text sequence $S$ and get the encoded representation $\textbf{H}^{C}$ through a transformer-based pre-trained language model. We can obtain the attention head matrix $\textbf{A}^{C}\in\mathbb{R}^{H\times l\times l}$ from the final transformer layer, where $H$ is the number of attention heads and $l$ represents the text sequence length. We then extract the contextual attention heads of the trigger and each argument through the pooling operation, denoted as $\textbf{A}^{C}_{t}\in\mathbb{R}^{l}$ and $\textbf{A}^{C}_{a_{i}}\in\mathbb{R}^{l}$. It’s important to note that since we do not know in advance what the argument is, the way it is obtained varies by method. For span-based methods, we consider the tokens within the span as the argument. For prompt-based methods, we treat the slot in the prompt template as the argument. We explain this in detail in Section 5.
Then for each argument, we get the contextual clue information $\textbf{c}_{a_{i}}\in\mathbb{R}^{d}$ by applying multiplication to the attentions and then normalizing the result:
$\centering\begin{array}[]{c}\textbf{p}^{C}_{a_{i}}=softmax(\textbf{A}^{C}_{a_{i}}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}_{a_{i}}={\mathbf{H}^{C}}^{T}\ \textbf{p}^{C}_{a_{i}},\end{array}\@add@centering$ (1)
where $\cdot$ means dot product and $\textbf{p}^{C}_{a_{i}}\in\mathbb{R}^{l_{w}}$ is the computed contextual attention weight vector. $T$ is the transpose symbol. $\textbf{c}_{a_{i}}$ is the enhanced context vector specific to the argument $a_{i}$ and trigger $t$. The detailed algorithm is shown in Algorithm 1.
Algorithm 1 Contextual Clues Aggregation
1:Input text sequence $S$, pre-trained transformer encoder $Encoder$
2:Context-enhanced argument representations $\textbf{c}_{a_{i}}$
3:$\textbf{H}^{C}\leftarrow Encoder(S)$triangleright Encode $S$ using the pre-trained transformer
4:Obtain attention head matrix $\textbf{A}^{C}\in\mathbb{R}^{H\times l\times l}$ from the final layer of $Encoder$
5:for each argument $a_{i}$ do
6:     Identify the index $t_{idx}$ of the trigger in the sequence
7:     Identify the index range $a_{i_{idx}}$ of the argument $a_{i}$ in the sequence
8:     Extract $\textbf{A}^{C}_{t}\in\mathbb{R}^{l}$ for the trigger by selecting the $t_{idx}$-th row from $\textbf{A}^{C}$
9:     Extract $\textbf{A}^{C}_{a_{i}}\in\mathbb{R}^{l}$ for the argument $a_{i}$ by aggregating rows in $a_{i_{idx}}$ from $\textbf{A}^{C}$
10:     Compute contextual attention weights: $\textbf{p}^{C}_{a_{i}}=\text{softmax}(\textbf{A}^{C}_{a_{i}}\cdot\textbf{A}_{t}^{C})$
11:     Obtain context vector: $\textbf{c}_{a_{i}}={\mathbf{H}^{C}}^{T}\textbf{p}^{C}_{a_{i}}$
12:end for
13:return $\textbf{c}_{a_{i}}$ for each argument $a_{i}$
4.2 Role-based Latent Information Guidance
In this section, we propose Role-based Latent Information Guidance (RLIG) module, consisting of Role-interactive Encoding and Role Information Fusion.
4.2.1 Role-interactive Encoding
In NLP applications, the inclusion of prefixes, like relation and other injected information, significantly enhances the performance of downstream tasks (Xu et al., 2022a; Hsu et al., 2023; Liu et al., 2023a). To effectively capture the semantic correlations between roles, we integrate the information of role types as prefixes into the input and utilize multi-head attention to facilitate interaction between the context and roles. We generate latent role embeddings by assigning unique special tokens   [^1] In our implementation, using [unused] tokens in BERT (Devlin et al., 2019) and introducing specific tokens for RoBERTa (Liu et al., 2019) to represent different roles in the framework of a pre-trained model, with each role type gaining a unique latent representation. Considering that role names carry valuable semantic information (Wang et al., 2022), we encase role names in special tokens indicative of their role types, selecting the embedding of the initial special token as the role’s representation. This approach ensures the capture of unique attributes and subtle semantic distinctions of each role. For instance, the role “Victim” is denoted as $[R_{0}]\quad Victim\quad[R_{0}]$, with $[R_{0}]$ representing the specialized token for “Victim”.
Algorithm 2 Role-based Latent Information Guidance
1:Input text sequence $S$, set of roles $\{r_{0},r_{1},\ldots,r_{n}\}$, pre-trained transformer encoder
2:Role-enhanced argument representations $\textbf{r}_{a_{i}}$
3:Enhance $S$ by concatenating role type elements:
4:$S^{\prime}=[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ ...\ w_{N}$
5:Encode $S^{\prime}$ using the pre-trained transformer to obtain $\mathbf{H}^{C}$ and $\mathbf{H}^{R}$
6:for each argument $a_{i}$ do
7:     Identify indices $t_{idx}$ and $a_{i_{idx}}$ for trigger $t$ and argument $a_{i}$ in $S^{\prime}$
8:     Extract $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ by selecting the attention vector corresponding to
9:  $t_{idx}$ from $\mathbf{H}^{R}$
10:     Extract $\textbf{A}^{R}_{a_{i}}\in\mathbb{R}^{l_{r}}$ by selecting the attention vector corresponding to
11:  $a_{i_{idx}}$ from $\mathbf{H}^{R}$
12:     Compute role attention weights: $\textbf{p}^{R}_{a_{i}}=\text{softmax}(\textbf{A}^{R}_{a_{i}}\cdot\textbf{A}^{R}_{t})$
13:     Obtain role-specific information: $\textbf{r}_{a_{i}}={\mathbf{H}^{R}}^{T}\textbf{p}^{R}_{a_{i}}$
14:end for
15:return $\textbf{r}_{a_{i}}$ for each argument $a_{i}$
Given the input sequence $S$, we enhance it by concatenating role type elements in the following manner:
$\quad S\ =\ \textsc{ $[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$}$
$\ \quad\quad\quad\quad w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N},$
where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the “none” category. By employing role-interactive encoding, we are able to get the context representation $\mathbf{H}^{C}\in\mathbb{R}^{l_{w}\times d}$, and the role representation $\mathbf{H}^{R}\in\mathbb{R}^{l_{r}\times d}$. Here, $l_{w}$ denotes the length of the list of word pieces, and $l_{r}$ represents the length of the list of roles.
4.2.2 Role Information Fusion
Through role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context. To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information. The aggregated role information $\textbf{r}_{a_{i}}\in\mathbb{R}^{d}$ for argument ${a_{i}}$ is obtained as follows:
$\begin{array}[]{c}\textbf{p}^{R}_{a_{i}}=softmax(\textbf{A}^{R}_{a_{i}}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}_{a_{i}}={\mathbf{H}^{R}}^{T}\ \textbf{p}^{R}_{a_{i}},\end{array}$ (2)
where $\textbf{A}^{R}_{a_{i}}\in\mathbb{R}^{l_{r}}$ represents the role attention for argument $a_{i}$, while $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ denotes the role attention associated with the trigger $t$. Additionally, the vector $\textbf{p}^{r}_{a_{i}}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles. $\textbf{r}_{a_{i}}$ is the aggregated role vector for argument $a_{i}$. The detailed algorithm is shown in Algorithm 2.
Then the enhanced context vector $\textbf{c}_{a_{i}}$ and role vector $\textbf{r}_{a_{i}}$ can be utilized to enhance the representation e of the argument $a_{i}$ to be predicted.
5 Adapting CARLG to Span-based and Prompt-based EAE Approaches
We instantiate the CARLG framework into two variants: the span-based $\text{CARLG}_{\text{span}}$ and the prompt-based $\text{CARLG}_{\text{prompt}}$, based on two types of current mainstream EAE approaches.
5.1 $\text{CARLG}_{\text{span}}$
We apply our CARLG framework to span-based EAE approaches, where the architecture is shown in Figure 4. Since the application of CARLG framework is consistent across span-based methods, we choose the architecture of TSAR (Xu et al., 2022b) as the foundation to illustrate the span-based baseline [^2] We extract the core architecture of TASR as the span-based baseline, removing the two-stream module and the AMR graph module., and we detail the application of our CARLG framework to other typical span-based methods in the experimental Section 6.2.
Figure 4: The main structure of $\text{CARLG}_{\text{span}}$. The input sequence, encompassing roles and event types, passes through a transformer-based encoder. This encoder generates output in the form of context representations, role representations, and attention heads. The Contextual Clue Aggregation (CCA) module dynamically consolidates contextual clues, while the Role-based Latent Information Generation (RLIG) module generates latent role embeddings and captures role correlations. Subsequently, the context vector is combined with the latent role vector to create the ultimate span representation. Finally, the classification module leverages this representation to forecast argument roles for all potential spans.
5.1.1 Role-interactive Encoder
We apply Algorithm 2 and construct a role-interactive encoder to capture role correlations. Following (Xu et al., 2022b; Ebner et al., 2020), we divide a document into several instances, where each instance only contains one event. For an instance, given the input document $\mathcal{D}$, the target event $e$ and its associated set of roles $\mathcal{R}_{e}=\{{r}_{1},{r}_{2},{r}_{3},...\}$, we enhance the input by concatenating these elements in the following manner:
$\quad S_{\text{span}}\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N}$
$\quad\quad\mathrm{[SEP]}\textsc{ \ $[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$},$
where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the “none” category. Next, we use a pre-trained transformer encoder to get the embedding of each token as follows:
$\mathbf{H}=\mathrm{Encoder}(S_{\text{span}}).$ (3)
@@ 5.1.1 Role-interactive Encoder
We can get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start $[E_{e}]$ and the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, where $l_{w}$ represents the length of the word pieces list. Notably, when the input sequence is longer than the maximum token length, we employ a dynamic window approach (Zhou et al., 2021; Zhang et al., 2021) to process the long sequence. We divide the sequence into overlapping windows and calculate the embeddings for each window. Next, we obtain the final representation by averaging the embeddings of tokens that overlap across various windows.
5.1.2 Contextual Clues Aggregation
We apply Algorithm 1 to the span-based representations, capturing context clues to enhance the representation of candidate spans. We employ the token-level attention heads $\textbf{A}^{C}\in\mathbb{R}^{H\times l_{w}\times l_{w}}$ obtained from the final transformer layer of the pre-trained language model. Here, $H$ represents the number of attention heads, and $l_{w}$ signifies the length of word pieces. Utilizing these attention heads, we compute the context attention $\textbf{A}^{C}_{i:j}\in\mathbb{R}^{l_{w}}$ for each candidate span spanning from $w_{i}$ to $w_{j}$, by applying average pooling:
$\textbf{A}^{C}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{C}_{h,m}.$ (4)
Then for span-trigger tuple $(s_{i:j},t)$, we get the contextual clue information $\textbf{c}_{s_{i:j}}\in\mathbb{R}^{d}$ which is crucial for the candidate span, by applying multiplication to the attentions and then normalizing the result:
$\begin{array}[]{c}\textbf{p}^{C}_{i:j}=softmax(\textbf{A}^{C}_{i:j}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}_{s_{i:j}}={\mathbf{H}^{C}}^{T}\ \textbf{p}^{C}_{i:j},\end{array}$ (5)
Figure 5: Detailed RLIG module for $\text{CARLG}_{\text{span}}$. Sub-figure (a) illustrates the information fusion and (b) illustrates the Latent Role Information Guidance operation.
5.1.3 Latent Role Information Guidance
Since span-based methods ultimately classify candidate spans based on the categories of roles, we propose Role Information Guidance on top of Algorithm 2 to better utilize the learned latent role representations, as shown in Figure 5.
Latent Role Information Fusion  Through the role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context. To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information. The aggregated role information $\textbf{r}_{s_{i:j}}\in\mathbb{R}^{d}$ for ${s_{i:j}}$ is obtained through contextual pooling, which involves adjusting the procedure described in Algorithm 2:
$\begin{array}[]{c}\textbf{A}^{R}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{R}_{h,m}\vspace{1.3ex},\\ \textbf{p}^{R}_{i:j}=softmax(\textbf{A}^{R}_{i:j}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}_{s_{i:j}}={\mathbf{H}^{R}}^{T}\ \textbf{p}^{r}_{i:j},\end{array}$ (6)
where $\textbf{A}^{R}\in\mathbb{R}^{H\times l_{w}\times l_{r}}$ represents the role attention heads obtained from the final transformer layer of the pre-trained model. $\textbf{A}^{R}_{i:j}\in\mathbb{R}^{l_{r}}$ represents the role attention for each candidate span, while $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ denotes the role attention associated with the trigger $t$. Additionally, the vector $\textbf{p}^{R}_{i:j}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles. We can fuse the contextual clue information $\textbf{c}_{s_{i:j}}$ and aggregated role information $\textbf{r}_{s_{i:j}}$into the average pooling representation as follows:
$\textbf{s}_{i:j}=tanh(\mathbf{W}^{Span}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{c}_{s_{i:j}};\textbf{r}_{s_{i:j}}]),$ (7)
where $\mathbf{W}^{Span}\in\mathbb{R}^{3d\times d}$ is learnable parameter. $\textbf{s}_{i:j}$ is the enhanced representation of the candidate span $s_{i:j}$.
Role Information Guidance  To maximize the utilization of the learned role representations, we introduce the Latent Role Information Guidance algorithm, tailored to the span-based architecture. We first map the original $d$-dimensional role features into a reduced dimension $d^{\prime}$ via a fully connected network to obtain the latent role information $\mathbf{H}^{r^{\prime}}\in\mathbb{R}^{l_{r}\times d^{{}^{\prime}}}$ as follows:
$\mathbf{H}^{r^{\prime}}=\mathbf{H}^{r}\mathbf{W}_{3}+\mathbf{b}_{w},$ (8)
where $\mathbf{W}_{3}\in\mathbb{R}^{d\times d^{{}^{\prime}}}$ and $\mathbf{b}_{w}\in\mathbb{R}^{d^{{}^{\prime}}}$ are learnable parameters. To fully utilize the expressive power of role embeddings, we propose a factorization-based alignment approach using Tucker decomposition  (Tucker et al., 1964). We define candidate span-trigger-role triples set $\mathcal{T}=\{<s_{i:j},r_{k},t_{e}>|s_{i:j}\in\mathcal{S},r_{k}\in\mathcal{R}_{e}\}$. Then we introduce a core tensor $\mathbf{Z}\in\mathbb{R}^{d_{i}\times d^{{}^{\prime}}}$ where $d_{i}$ is the hidden dimension of the combined representation $\mathbf{I}_{i:j}$. $\mathbf{I}_{i:j}=FFN([\mathbf{h}_{t};\textbf{s}_{i:j}])$, and $FFN$ is the full-connected neural network. For simplity, the validity for each triple $<s_{i:j},r_{k},t_{e}>\in\mathcal{T}$ is scored as follows:
where ${\widehat{\mathbf{r}}}_{\mathrm{k}}=\mathbf{H}^{r^{\prime}}_{k,:}$ represents the transformed role features and ×$\times_{n}$ denotes the tensor product operation along the $n^{th}$ dimension. The calculation of $\varphi$ for all triples is performed concurrently via a batched tensor product technique, and the model is trained employing focal loss as described in reference (Lin et al., 2017).